# Hybrid Monte Carlo for Failure Probability Estimation with Gaussian Process Surrogates

### v1

@@ 5 RAE-2822 airfoil computer experiment
The RAE-2822 airfoil computer experiment simulates the flow of air around an aircraft wing. The simulation suite is solved via SU2, and the software is publicly available (Economon et al.,, 2016). The experiment accepts seven inputs: four shape parameters ($S_{i}$ for $i=1,\dots,4$) and 3 environmental parameters, namely angle of attack, Reynolds number, and Mach number. We are interested in the lift over drag ($L/D$) ratio. $L/D$ values that are too low represent inefficient flight conditions. We define a failure threshold at $t=3$ such that $L/D<3$ indicates failure. [Note the change from $f(\mathbf{x})>t$ to $f(\mathbf{x})<t$ does not affect the implementation of our method.] This computer experiment requires nearly 30 minutes of compute time with serial execution, but reductions are possible with parallel execution.
@@
$S_{i}$ $\sim\mathrm{Unif}(-1e-4,1e-4)\;\;\textrm{for}\;\;i=1,\dots,4$
@@
Reynolds number $\sim\mathcal{N}\left(\mu=1e7,\sigma=1e6\right)\;\;\textrm{truncated to}\;\;[5e6,1.5e7]$
@@
We fix our total budget at $T=500$ simulator evaluations. Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment. We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package. We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds. Given the input dimension, we require atleast 40 observed failures before we consider halting surrogate training. Since the true failure probability is unknown, we repeat this entire exercise with a newly randomized starting design for some additional context. Figure 6 reports the progress in $\hat{\alpha}$ over the course of each design.
Figure 6: Two surrogate MC failure probabilty estimates for the airfoil computer experiment, updated every 10th acquisition with final estimates in dashed red. Grey shading indicates $\hat{\alpha}\pm\hat{\sigma}_{\alpha}$ for the final estimates.
Our stopping criterion halted the sequential designs at $N=180$ and $N=160$. For each repetition, we collect a separate random sample of size $M=2.5\times 10^{6}$, then identify the remaining points with the highest entropy, and evaluate the true simulator at these locations. After updating the DGP surrogate, we obtain the final hybrid Monte Carlo estimates of $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$, respectively. These values are indicated by the red dashed lines in Figure 6. The grey shading shows $\hat{\alpha}\pm\hat{\sigma}_{\alpha}$, calculated from these final hybrid estimates. Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$. Additionally, notice how the final hybrid estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had estimated, closer to the estimate of our first attempt (left panel). We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
@@ 6 Discussion
We presented a hybrid Monte Carlo method for estimating failure probabilities of expensive computer experiments. Our hybrid method leverages the uncertainty quantification provided by a surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty. We limited our experiments to GP and DGP surrogates, but our proposed estimator is applicable to any surrogate which provides posterior predictive distributions (required for calculation of entropy). Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
@@
Our hybrid MC estimator is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate. Posterior predictions must be gathered at all $M$ sample points. While surrogate predictions are very cheap, collecting millions of them is not necessarily trivial. Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason. We could tackle smaller failure probabilities, but it would require larger $M$, which would eventually become impractical. Our approach may simply not be suited to very rare failure probabilities.
The efficacy of any surrogate-informed failure probability estimation relies on the surrogate’s ability to accurately classify failures. Although our hybrid estimator is generally applicable, we strongly recommend the use of contour locating sequential designs to train the surrogate. We proposed a budgeting scheme which informs allocation of simulator evaluations between surrogate training and failure probability estimation based on progress in the surrogate Monte Carlo estimate. If a budget is severly restricted, the budget could be exhausted training the surrogate before our proposed stopping criterion is met. While this is not ideal, we contend that fully-training a surrogate and using a surrogate MC estimate is the best option in this scenario. Particularly in light of the data hungry nature of IS-based methods, which perform poorly in small data settings.
@@
When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024). There is arguably no need to train a surrogate in a region where inputs will not plausibly occur, such as in the outer corners of the Herbie function surface (Figure 1, center). Incorporating this information in the contour location stage could improve failure probability estimation downstream. But in the case of rare probabilities, it may be hard to draw the line between regions that are plausible and ones that are not. This avenue is currently underexplored.
@@
Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 3.2. In each case, failures are defined as $f(\mathbf{x})>t$.
@@
where $\Phi$ is the standard Gaussian CDF. We use $d=4$ and set the failure threshold at $t=0$. We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.5,0.085)$ for $i=1,\dots,4$, truncated to $[0,1]$.

### v2

@@ 5 RAE-2822 airfoil computer experiment
The RAE-2822 airfoil computer experiment simulates the flow of air around an aircraft wing. The simulation suite is solved via SU2, and the software is publicly available (Economon et al.,, 2016). We consider seven inputs: four shape parameters ($S_{i}$ for $i=1,\dots,4$) and 3 environmental parameters, namely angle of attack, Reynolds number, and Mach number. We are interested in the lift over drag ($L/D$) ratio. $L/D$ values that are too low represent inefficient aircraft performance. We define a failure threshold at $t=3$ such that $L/D<3$ indicates failure. [Note the change from $f(\mathbf{x})>t$ to $f(\mathbf{x})<t$ does not affect the implementation of our method.] This computer experiment requires nearly 30 minutes of compute time with serial execution, but reductions are possible with parallel execution.
@@
$S_{i}$ $\sim\mathrm{Unif}(-1\times 10^{-4},1\times 10^{-4})\;\;\textrm{for}\;\;i=1,\dots,4$
@@
Reynolds number $\sim\mathcal{N}\left(\mu=1\times 10^{7},\sigma=1\times 10^{6}\right)\;\;\textrm{truncated to}\;\;[5\times 10^{6},1.5\times 10^{7}]$
@@
We fix our total budget at $N=500$ simulator evaluations and use $M=2.5\times 10^{6}$ for the size of $X_{M}$. Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment. We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package. We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds. Once we halt contour location (Section 3.1), we use entropy to inform our hybrid MC estimator (Section 3.2). Given the larger input dimension, we require at least 40 observed failures before we consider halting surrogate training. Since the true failure probability is unknown, we opt to repeat this entire exercise twice, with a newly randomized starting design. We consider the agreement between the two randomized repetitions as an indication that our procedure is well-calibrated.
Figure 7: Progress in $\hat{\alpha}$ across two contour locating designs for the airfoil computer experiment. Proposed two-stage failure probability estimate shown in dashed red. Grey shading indicates $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$ for these final estimates. Left/right panels show two repetitions with newly randomized starting designs.
Figure 7 reports the progress in $\hat{\alpha}$ over the course of each design. Our stopping criterion halted the sequential designs at $N=180$ and $N=160$, leaving $B=320$ and $B=340$ for the second design stage, respectively. Our two hybrid Monte Carlo estimates were $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$. These values are indicated by the red dashed lines in Figure 7. The grey shading shows $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$, calculated from these final estimates. Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$. Additionally, notice how the final hybrid MC estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had originally estimated, closer to the estimate of our first attempt (left panel). We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
@@ 6 Discussion
We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments. Starting with a contour locating sequential design, we halt contour location once surrogate training has plateaued, as indicated by practical convergence of a surrogate Monte Carlo estimator. Then we use the remaining budget of simulator evaluations on the Monte Carlo samples with highest entropy, to ultimately return a hybrid MC estimate of $\alpha$. Our design strategically leverages the uncertainty quantification provided by the surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty. While we limited our experiments to GP and DGP surrogates, our proposed scheme is applicable to any surrogate which provides posterior predictive distributions (which are required for entropy calculations). Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
@@
Our two-stage design is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate. Posterior predictions must be computed at all $M$ samples multiple times. While surrogate predictions are very cheap, collecting billions of them is not necessarily trivial. Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason. We could tackle smaller probabilities, but it would require larger $M$, which would eventually become impractical. Although we are focused on expensive simulators whose computational costs overshadow the cost of these surrogate predictions, our approach may simply not be suited to extremely rare failure probabilities.
In all of our examples, our stopping criterion terminated contour location before the complete budget ($N$) was exhausted. If a simulation budget is severely restricted, the budget could be exhausted before our proposed stopping criterion is met. While we prefer to have additional evaluations reserved for the estimation stage, we contend that using the entire budget for contour location to provide a surrogate MC estimate is the best option in this scenario. As we’ve shown, importance sampling based methods perform poorly when data is limited.
We have exclusively focused on deterministic computer experiments. In fact, the noise-free nature of the simulator is what enabled us to frame our contribution as a two-stage design. Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$. Thus our hybrid MC estimator (7) is really a surrogate MC estimator (4), with a strategically trained surrogate. Different design choices could be warranted for stochastic simulators (Baker et al.,, 2022). In particular, the hybrid estimator of Eq. (7) should not make use of noisy $f(\mathbf{x})$. Replicates in the design may be crucial (Binois et al.,, 2019).
@@
When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024). There is arguably no need to train a surrogate in a region where inputs will never occur. Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as the black open circles around the edges of Figure 3). We suspect there is a delicate balance that would work well here, but this avenue is currently underexplored.
Estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization. For example, in the context of our airfoil experiment, we may wish to find the optimal wing configuration given that the probability of failure remains below a specified threshold. The ability to effectively estimate failure probabilites from limited data, which we have tackled here, is integral to the success of this larger enterprise.
@@
Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 4. In each case, failures are defined as $f(\mathbf{x})>t$.
@@
where $\Phi$ is the standard Gaussian CDF. We use $d=4$ and set the failure threshold at $t=0$. We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.6,0.11)$ for $i=1,\dots,4$, truncated to $[0,1]$.